### v4

@@ 4.2 Model Comparison and Evaluation
Table 4: Summary of the trainable parameters and key outputs for the selected FNN and QML models, including the best average RMSE, the epoch at which the lowest average RMSE is achieved, the average standard deviation of the RMSE, and the average execution time per epoch. All values are calculated over four cross-validation folds with one hundred epochs.
@@
The results, summarized in Figure 8 and Tables 4, provide several key insights. First, the QML models (using Amplitude Encoding and Angle Encoding) outperform the classical FNN model, showcasing superior generalizability. They achieve lower average RMSE on test data with fewer epochs (see columns two and three of Table 4). Moreover, their stability is evidenced by the smaller standard deviation observed across the four cross-validation folds (column four of Table 4). These findings underscore the effectiveness of replacing a classical layer with a parameterized quantum circuit (PQC) in reducing overfitting, particularly in tasks such as recovery rate prediction.
@@
Second, the QML model with Amplitude Encoding (QML Amp) demonstrates the best overall performance. Despite its simplicity and minimal number of trainable parameters, it achieves the lowest average RMSE, requires the fewest epochs, and exhibits the smallest average standard deviation.
@@
When comparing the QML models, the Amplitude Encoding model shows more remarkable advantages over Angle Encoding in both stability and accuracy, as detailed in Table 4 and Figure 8. While both models outperform the classical FNN with a comparable number of trainable parameters, the Amplitude Encoding model consistently achieves lower RMSE values and smaller standard deviations across multiple runs. This superior performance highlights the effectiveness of Amplitude Encoding in capturing and utilizing input information more efficiently. Its enhanced stability and accuracy can be attributed to its ability to represent input data compactly without requiring additional auxiliary layers to reduce input dimensions. By avoiding this added complexity, the Amplitude Encoding model minimizes the risk of overfitting, particularly with limited data. In contrast, the Angle Encoding model relies on an auxiliary classical layer to align the input size with the number of PQC qubits, which can introduce unnecessary complexity and negatively impact performance. Furthermore, the compactness of Amplitude Encoding enables it to encode a larger amount of information into the quantum state using fewer qubits, giving it a significant advantage over Angle Encoding, which scales less efficiently. These attributes make Amplitude Encoding a compelling choice for tasks demanding high accuracy and strong generalization, as reflected in our results.
@@ 5 Conclusion and Discussion
In this work, we proposed and evaluated a quantum machine learning (QML) model with Amplitude Encoding for recovery rate prediction, comparing its performance with a QML model using Angle Encoding and a classical feedforward neural network (FNN). Our experiments demonstrate that the QML model with Amplitude Encoding outperforms both the QML model with Angle Encoding and the classical FNN in terms of accuracy, stability, and generalization. The Amplitude Encoding model’s lower RMSE and smaller standard deviation across multiple training runs highlight its superior ability to avoid overfitting, making it a promising approach for prediction tasks involving complex data.
@@
Although the QML model with Angle Encoding shows improvements over the classical FNN, this is not as significant as the one achieved by the QML with Amplitude encoding. The performance of the QML Ang is hindered by the added complexity of the auxiliary classical layer, which increases the risk of overfitting. In contrast, the simplicity of the Amplitude Encoding model, with fewer layers and parameters, allows it to achieve better results with a more stable training process.
From a computational perspective, our results indicate that the QML model with Amplitude Encoding is slightly more time-efficient than the Angle Encoding model in simulation. However, the practical implementation of Amplitude Encoding on real quantum hardware faces challenges due to the need for deeper quantum circuits, which are more susceptible to noise and decoherence. These factors highlight the importance of ongoing advancements in quantum hardware to fully leverage the potential of Amplitude Encoding in practical applications. Despite these challenges, the superior simulation performance and scalability of Amplitude Encoding underscore its promise as a robust approach for quantum machine learning tasks.
Overall, our study provides valuable insights into the potential of quantum-enhanced machine learning and demonstrates the advantages of Amplitude Encoding for certain prediction tasks. In future work, we aim to explore two key directions. First, to improve the scalability and robustness of quantum models, adding more qubits in QML with Amplitude Encoding can enable the handling of exponentially higher-dimensional datasets. However, when the dataset dimension is not a perfect power of two, challenges such as data padding with appropriate schemes must be addressed. Second, regarding the practical deployment of quantum hardware, while it has the potential to achieve more time-efficient training compared to classical models, it is crucial to study the impact of noise and decoherence. Understanding and potentially leveraging these phenomena could lead to more reliable and efficient training processes.
@@ Acknowledgements
Appendix A The Classical FNNs and overfitting
@@
$n$ Hidden ##\## Parameters Best Average RMSE # Average STD
8 67,857 0.246 0.018
@@
Appendix B Performance increasing the number of qubits in the QML model with Angle Encoding
In the architecture proposed in Section 3, the number of qubits can be adjusted to identify the optimal configuration for the QML model using Angle Encoding. Figure 10 presents the average RMSE and standard deviation on the test data for two configurations: one with six qubits and another with fourteen qubits in the PQC. As in the previous section, Table 6 summarizes the best RMSE and corresponding standard deviation values observed across ten experiments for different qubit configurations.
Figure 10: The average test RMSE and STD calculated over four corss-validation folds of the QML with Angle Encoding for seven and fourteen qubits in the PQC.
$n$ Qubits ##\## Parameters Best Average RMSE Average STD
6 67,353 0.241 0.012
@@ Appendix B Performance increasing the number of qubits in the QML model with Angle Encoding
From Table 6 and Figure 10, we can draw two key conclusions. First, the QML model with Angle Encoding does not demonstrate improved predictive performance with an increasing number of qubits. While there is a slight improvement with six and seven qubits, this effect saturates after eight qubits. Second, a comparison of Tables 5 and 6 reveals that the QML model with Angle Encoding offers a slight improvement over the classical FNN. In all reported cases, the QML Ang outperforms the classical FNN model in both stability (lower standard deviation) and effectiveness (lower average RMSE).
@@
As a final remark, Figure 11 shows that the average execution time per epoch of the QML model with Angle Encoding increases exponentially with the number of qubits, consistent with the resource demands of the state-vector default.qubit simulator. For illustration purposes, in the same Figure 11, we also mark the execution time of the QML with Amplitude Encoding proposed in the current work, highlighting the slight advantage performance.
Figure 11: Average execution time per epoch in the QML with angle encoding (blue line) and amplitude encoding (orange marker).

### v5

@@ 5.3.1 Cross and Leave-One-Out-Cross Validation
Table 7: Summary of the trainable parameters and key outputs for the selected FNN and QML models, including the best average RMSE, the epoch at which the lowest average RMSE is achieved, the average standard deviation of the RMSE, and the average execution time per epoch. All values are calculated over four cross-validation (CV) folds and leave-one-out (LOOCV), with one hundred epochs.
@@
The results, summarized in Figure 7(a)7(b) and Tables 7, provide several key insights. First, the QML models (using Amplitude Encoding and Angle Encoding) outperform the classical FNN model, showcasing superior generalizability. They achieve lower average RMSE in both cross and leave-one-out cross validation, with fewer epochs (see columns three and four of Table 7). Moreover, their stability is evidenced by the smaller standard deviation observed across the four cross-validation folds and all left-out observations (column five of Table 7). These findings underscore the effectiveness of replacing a classical layer with a QNN in reducing overfitting, particularly in tasks such as recovery rate prediction.
@@
Second, the QML model with Amplitude Encoding demonstrates the best overall performance. Despite its simplicity and minimal number of trainable parameters, it achieved the lowest average RMSE, requiring the fewest epochs, with the smallest average standard deviation.
5.3.2 Statistical significance
Figure 8: A summary of test significance across each fold and each epoch. The green (red) boxes indicate that the $DM$ is negative (positive), which means that the model (a) performs better (worse) than model (b). The intensity of the color indicates the p-values: when closer to 0.05, white, which is less significant, and greener (redder) when closer to zero, which indicates higher significance.
Figure 8 reports the p-values for the hypothesis test $DM_{a,b}=0$. The color intensity increases as the p-value approaches zero, indicating a stronger statistical difference between the two models. Boxes are shown in grey when the p-value exceeds the critical threshold of 0.05, suggesting no significant difference between the models. When $DM<0$ (respectively, $DM>0$), boxes are colored green (respectively, red), indicating that model $a$ performs significantly better (worse) than model $b$.
Pairwise comparisons of the selected models show that the differences between the QML with Amplitude Encoding and the FNN, and between the QML with Amplitude Encoding and QML with Angle Encoding, are significantly greater than zero across nearly all folds and training epochs (regardless of the starting points in training). This confirms that the QML with Amplitude Encoding model statistically outperforms both the FNN and QML with Angle Encoding models. However, the comparison between FNN and QML with Angle Encoding does not yield consistent results, as the significance tests frequently fail throughout the training process.
5.3.3 Discussion
@@
When comparing the QML models, the Amplitude Encoding model shows more remarkable advantages over Angle Encoding in both stability and accuracy, as detailed in Table 7, and Figures 78. While both models outperform the classical FNN with a comparable number of trainable parameters in terms of average RMSE, the Amplitude Encoding model consistently achieves lower RMSE values and smaller standard deviations across multiple experiments. This superior performance highlights the effectiveness of Amplitude Encoding in capturing and utilizing input information more efficiently. Its enhanced stability and accuracy can be attributed to its ability to represent input data compactly without requiring additional auxiliary layers to reduce input dimensions. By avoiding this added complexity, the Amplitude Encoding model minimizes the risk of overfitting, particularly with limited data. In contrast, the Angle Encoding model relies on an auxiliary classical layer to align the input size with the number of PQC qubits, which can introduce unnecessary complexity and negatively impact performance. Furthermore, the compactness of Amplitude Encoding enables it to encode a larger amount of information into the quantum state using fewer qubits, giving it a significant advantage over Angle Encoding, which scales less efficiently. These attributes make Amplitude Encoding a compelling choice for tasks demanding high accuracy and strong generalization, as reflected in our results.
@@
5.4 Comparison evaluation with XGBoost
5.4.1 Cross and Leave-One-Out-Cross Validation
As in the previous section, Figures 9(a) and 9(b) show the average RMSE (solid lines) and the corresponding standard deviations (shaded areas) obtained by the QML with Amplitude Encoding and XGBoost models under four-fold and leave-one-out cross-validation, respectively. Table 8 reports the best average test RMSE achieved during training, the associated average RMSE standard deviation, and the average execution time per epoch (boosted tree optimization for XGBoost).
(a) Four-fold cross validation (CV)
(b) Leave-one-out cross validation (LOOCV)
Figure 9: Average test RMSE (solid lines) and standard deviation (shaded area) calculated on a four-fold cross validation (CV left panel) and leave-one-out cross validation (LOOCV right panel) for the XGBoost and QML with Amplitude Encoding. For the sake of visualization, given the high variance of the RMSE calculated on each left-out observation, the shaded area on the right panel indicates a scaled variance by a factor of four.
Model # Parameters Avg. Time per Epoch (s) Best Avg. RMSE At Epoch/Tree # Avg. STD
XGBoost 6,400 0.02 0.222 (CV) 100 (CV) 0.010 (CV)
0.141 (LOOCV) 100 (LOOCV) 0.039 (LOOCV)
QML Amplitude 65,825 0.73 0.228 (CV) 55 (CV) 0.008 (CV)
Encoding 0.136 (LOOCV) 86 (LOOCV) 0.043 (LOOCV)
Table 8: Summary of the trainable parameters and key outputs for the selected XGBoost and QML with Amplitude Encoding models, including the best average RMSE, the epoch at which the lowest average RMSE is achieved, the average standard deviation of the RMSE, and the average execution time per epoch. All values are calculated over four cross-validation (CV) folds and leave-one-out (LOOCV), with one hundred epochs (trees).
In this case, as shown in Figure 9(a) and Table 8, the XGBoost regression model achieves a lower RMSE value on the four-fold cross-validation setting. However, as we can see from column four of Table 8, to achieve the convergence XGBoost model needs multiple trees, potentially compromising the performance of the model (a deeper discussion about the scalability and the convergence of XGBoost is provided in Appendix A.3). Additionally, as illustrated in Figure 9(b) and column three of Table 8 in the leave-one-out validation, after one hundred boosted trees, the average RMSE is still higher than the one achieved by our proposed QML model.
5.4.2 Statistical significance
Figure 10: A summary of test significance across each fold and each epoch. The green (red) boxes indicate that the $DM$ is negative (positive), which means that the model (a) performs better (worse) than model (b). The intensity of the color indicates the p-values: when closer to 0.05, white, which is less significant, and greener (redder) when closer to zero, which indicates higher significance.
As in the previous section, Figure 10 highlights the instances where the QML model with Amplitude Encoding significantly outperforms the XGBoost model. Although XGBoost achieved lower RMSE values, its higher variance prevents it from being statistically significantly better than the QML model with Amplitude Encoding across most epochs and folds.
5.4.3 Discussion
The comparative analysis of XGBoost and QML with Amplitude Encoding models reveals a nuanced performance landscape. As shown in Figure 9(a) and Table 8, in standard cross-validation, XGBoost achieved a lower average RMSE compared to QML, though this came at the cost of higher variance, which increases during training. Conversely, under leave-one-out cross-validation, the QML model slightly outperforms XGBoost in terms of RMSE.
When examining performance across training epochs for QML and the sequential addition of trees in XGBoost, no statistically significant advantage consistently emerged for XGBoost over the proposed QML model, as pictured in Figure 10. Notably, in the early stages of training, QML demonstrated a statistically significant advantage over XGBoost, suggesting faster initial learning.
Column three of Table 8 shows that, while XGBoost exhibited faster execution times, this comparison is not meaningful: XGBoost was executed on GPU hardware, whereas QML relied on classical quantum simulation. Moreover, as illustrated in column 8, although XGBoost typically requires fewer tunable parameters per tree, its convergence depends on a large number of boosted trees. This iterative process increases the overall model complexity and the effective number of parameters (further discussion about the scalability of XGBoost is reported in Appendix A.3).
Taken together, these findings suggest that, despite current hardware limitations, QML remains a viable and competitive approach for recovery rate prediction. With further in-depth research, fine-tuning, and architectural adaptation, its complexity could potentially be reduced and its performance enhanced.
5.5 Further analysis about performance differences across Neural Network models
In this section, we highlight key architectural and training-related features of the proposed QML model with Amplitude Encoding that contribute to its competitive performance compared to both classical baselines and QML models using Angle Encoding.
First, Amplitude Encoding allows the QNN to represent all input features simultaneously in a compact quantum state. Unlike Angle Encoding—which encodes each feature individually via rotation gates and often requires dimensionality reduction to fit within qubit number constraints—Amplitude Encoding enables the representation of $2^{n}$ features using only $n$ qubits. This efficiency eliminates the need for auxiliary preprocessing steps and preserves more information, which is particularly beneficial for high-dimensional, small-sample tasks.
Second, the QML with Amplitude Encoding introduces a Parameterized Quantum Circuit (PQC) as a unitary transformation within the neural architecture. This unitary layer contributes to better generalization, reducing the risk of overfitting [14, 60]. Unlike classical Feedforward Neural Network (FNN), which often requires more complex architectures and additional constraints to prevent overfitting, the QNN benefits from a more compact yet expressive representation.
Third, the number of trainable parameters in the amplitude-encoded QNN scales logarithmically with the number of input features. This is in contrast to classical FNNs and angle-encoded QNN, where parameter count typically scales linearly with input size. The resulting architecture is not only more compact but also exhibits more stable training dynamics.
Together, these factors explain the superior performance of the QML model with Amplitude Encoding in both predictive accuracy and robustness.
@@ 6 Conclusion and Discussion
In this work, we proposed and evaluated a QML model with Amplitude Encoding for recovery rate prediction, comparing its performance with a QML model using Angle Encoding, classical Feedforward Neural Network (FNN), and XGBoost.
Our experimental results show that the QML model with Amplitude Encoding consistently outperforms the QML model with Angle Encoding and classical FNN in terms of accuracy, stability, and generalization. Its lower RMSE, statistically significant reduction in prediction error, and smaller standard deviation across multiple training runs and cross-validation settings highlight its robustness and reduced tendency to overfit, making it a promising candidate for complex prediction tasks.
@@
Although the QML model with Angle Encoding shows improvements over the classical FNN, this is not as significant as the one achieved by the QML with Amplitude encoding. The performance of the QML with Angle Encoding is hindered by the added complexity of the auxiliary classical layer, which increases the risk of overfitting. In contrast, the simplicity of the Amplitude Encoding model, with fewer layers and parameters, allows it to achieve better results with a more stable training process.
When compared to XGBoost, the proposed QML model with Amplitude Encoding does not demonstrate clearly superior performance. In fact, it yields a higher average RMSE in cross-validation settings, albeit with a smaller standard deviation. However, due to the higher variance observed in XGBoost, its performance improvement is not statistically significant. Moreover, in the leave-one-out cross-validation, the proposed QML model outperforms XGBoost. These findings suggest that further tuning of the QML model’s parameters and architecture could potentially enhance its performance and make it more competitive.
From a computational perspective, our results indicate that the QML model with Amplitude Encoding is slightly more time-efficient in simulation than the model using Angle Encoding. This is expected, as the Angle Encoding model requires an additional auxiliary layer in its architecture. Comparing the time performance of the QML model with the other classical models is not meaningful at this stage. A fair assessment of time efficiency can only be made when the QML model is implemented on real quantum hardware. That said, the practical implementation of Amplitude Encoding on current quantum devices remains challenging due to the need for deeper quantum circuits (see next Section 6.1), which are more susceptible to noise and decoherence.
6.1 Limitation
While this study demonstrates the potential of QML models in recovery rate modeling, several limitations remain, particularly in relation to hardware implementation, noise handling, feature encoding strategies, and overall Machine Learning architecture.
First, all QML models in this paper are developed and tested using quantum simulators rather than real quantum hardware. Simulation allows for noiseless training and evaluation, which does not fully reflect the limitations of existing quantum devices. In practice, current quantum processors suffer from limited qubit counts, gate fidelity issues, and short coherence times. As a result, implementing QNNs on real hardware remains a significant challenge. For example, Angle Encoding strategies typically require more qubits but produce shallower circuits. At the same time, Amplitude Encoding is more qubit-efficient but results in deeper and more complex circuits that are more sensitive to noise [67, 86, 95].
Second, our current implementation does not explicitly model or mitigate noise effects such as decoherence and gate errors. These sources of quantum noise can significantly deteriorate model performance when deploying on real quantum hardware. Techniques such as noise-aware training, error mitigation, and quantum error correction are potential directions for future work, especially as quantum hardware matures [97, 78]. However, a preliminary evaluation of the proposed models on real quantum hardware and noisy classical simulators is provided in Appendix B. Although these results are still in the early stages and do not yet support strong conclusions, they indicate that the proposed models may offer sufficient noise resilience and retain promising performance even under realistic noise conditions.
Third, the encoding of high-dimensional features into quantum circuits presents practical challenges. Amplitude Encoding requires the feature vector to be of length $2^{n}$, necessitating the use of padding or zero-filling techniques when the number of features does not match a power of two [87]. While padding allows compatibility, it may introduce redundancy or distort input distributions. Alternatively, Angle Encoding avoids this restriction but becomes impractical when the number of features is very large, as it requires one qubit per feature. In such cases, dimensionality reduction techniques or hybrid classical-quantum approaches are needed to reduce the input space before quantum processing.
In addition to the input formatting challenges, the number of qubits required for Amplitude Encoding still grows logarithmically with the number of input features. While this scaling is significantly more efficient than linear methods such as Angle Encoding, it can still become a bottleneck in practice, given the limited number of qubits currently available on Noisy Intermediate-Scale Quantum (NISQ) devices. For example, encoding 1,024 features would require at least 10 qubits, which does not account for additional qubits needed for variational processing or error mitigation. As a result, even moderate feature dimensions may exceed the capabilities of current hardware. This constraint highlights the need for preprocessing strategies such as dimensionality reduction or feature grouping prior to quantum encoding, or the development of more hardware-efficient encoding methods.
Moreover, the high dimensionality of the input is initially processed by a classical master layer, which may reduce the quantum speedup benefits by introducing computational bottlenecks. This reflects a broader trade-off in hybrid quantum-classical architectures. While the current design represents one viable approach, it is not necessarily optimal.
Finally, like classical neural networks, our QML model inherits challenges such as reduced interpretability and increased computational cost during training and inference on high-dimensional datasets [45, 35].
Overall, these limitations highlight both the promise and the current barriers to deploying QML in real-world financial applications. Overcoming these challenges will be essential as quantum technologies advance toward practical deployment.
6.2 Future Works
Building on the current findings, several directions for future research emerge, both from a methodological and an applied financial perspective.
First, future work can address the limitations discussed above by exploring more robust and efficient quantum model implementations. In particular, alternative encoding strategies such as basis encoding, hybrid schemes [83], or data re-uploading [76] could be evaluated for their effectiveness and scalability. When using Amplitude Encoding, padding strategies for non-power-of-two feature dimensions could be systematically tested to assess their impact on model accuracy and stability. Similarly, further exploration of dimensionality reduction techniques would be valuable when using Angle Encoding with high-dimensional inputs. From a hardware perspective, the deployment of QNNs on real quantum devices will require advances in circuit optimization, efficient gradient calculation (e.g., parameter-shift rules [61], adjoint differentiation [46]), and noise mitigation strategies to ensure reliable training and inference. These developments are essential for assessing the practical viability of QML models in financial applications.
Second, the economic implications of improved recovery rate prediction warrant further investigation. A natural extension would be to evaluate whether the QML-based recovery forecasts can inform profitable trading strategies. For instance, one could simulate a strategy that selectively purchases defaulted bonds with high predicted recovery rates and avoids or shorts those with low expected recoveries. However, the feasibility and profitability of such strategies depend on market efficiency, liquidity, and transaction costs. It is possible that bond prices already reflect recovery expectations due to market participants’ access to similar information. Therefore, future work should explore whether the QML model provides incremental economic value beyond existing credit risk models and whether its predictions offer statistically and economically significant alpha.
Third, an important direction for future research is to evaluate the model’s adaptability to dynamic financial environments. In our current study, the rarity of default events and limited sample size make it difficult to assess temporal generalization directly. However, as more recovery data become available or as higher-frequency credit risk datasets (e.g., CDS spreads, market-based distress indicators) are integrated, future studies could investigate whether QML models maintain predictive performance across different market regimes or macroeconomic cycles. One possible approach is to incorporate time-varying covariates or macro factors into the model architecture, or to test model transferability across temporal subsamples. Additionally, evaluating the robustness of QML under concept drift or structural shifts in default behavior would help assess their practical viability in real-world, evolving financial systems.
Together, these directions highlight the dual potential of QML for both technical innovation and practical impact in credit risk modeling.
@@ Acknowledgements
Appendix A Scalability of alternative regression models
In this appendix, we examine how the alternative models—FNN, QML with Angle Encoding, and XGBoost—respond to increasing architectural complexity. Specifically, we analyze the impact of increasing the number of hidden nodes in the FNN, the number of qubits in the QML model, and both the depth and number of trees in XGBoost.
A.1 The Classical FNNs and overfitting
@@
Model $n$ Hidden ##\## Parameters Best Avg. RMSE # Avg. STD
QML Angle Encoding 8 67,857 0.246 0.018
@@
A.2 Performance increasing the number of qubits in the QML model with Angle Encoding
In the architecture proposed in Section 4, the number of qubits can be adjusted to identify the optimal configuration for the QML model using Angle Encoding. Figure 12(a) presents the average RMSE and standard deviation on the test data for two configurations: one with six qubits and another with fourteen qubits in the PQC. As in the previous section, Table 10 summarizes the best RMSE and corresponding standard deviation values observed across ten experiments for different qubit configurations.
(a) Test RMSE
(b) Average Time per epoch
Figure 12: (Left panel) The average test RMSE and STD calculated over four corss-validation folds of the QML with Angle Encoding with seven and fourteen qubits in the PQC. (Right Panel) Average execution time per epoch in the QML with Angle Encoding (blue line) and Amplitude Encoding (orange marker).
Model $n$ Qubits ##\## Parameters Best Avg. RMSE Avg. STD
QML Angle Encoding 6 67,353 0.241 0.012
@@ A.2 Performance increasing the number of qubits in the QML model with Angle Encoding
From Table 10 and Figure 12(a), we can draw two key conclusions. First, the QML model with Angle Encoding does not demonstrate improved predictive performance with an increasing number of qubits. While there is a slight improvement with six and seven qubits, this effect saturates after eight qubits. Second, a comparison of Tables 9 and 10 reveals that the QML model with Angle Encoding offers a slight improvement over the classical FNN. In all reported cases, the QML with Angle Encoding outperforms the classical FNN model in both stability (lower standard deviation) and effectiveness (lower average RMSE).
@@
As a final remark, Figure 12(b) shows that the average execution time per epoch of the QML model with Angle Encoding increases exponentially with the number of qubits, consistent with the resource demands of the state-vector default.qubit simulator. For illustration purposes, in the same Figure 12(b), we also mark the execution time of the QML with Amplitude Encoding proposed in the current work, highlighting the slight advantage performance.
A.3 Performance increasing the number of trees in XGBoost
This Appendix illustrates a benchmark using XGBoost on the same regression task. We systematically varied the maximum tree depth parameter and recorded the corresponding test RMSE on a four-fold cross-validation, and the number of trees required to converge (Table 11 and Figure 13).
Model Max Tree depth # Parameters Best Avg. RMSE At Epoch/ Tree # Avg. STD
XGBoost 4 $\approx 8,000$ 0.223 495 0.011
5 $\approx 10,000$ 0.223 305 0.015
6 $\approx 18,000$ 0.220 270 0.011
7 $\approx 22,000$ 0.222 170 0.014
8 $\approx 18,000$ 0.224 70 0.014
9 $\approx 23,000$ 0.227 45 0.014
10 $\approx 40,000$ 0.230 43 0.015
11 $\approx 81,000$ 0.231 42 0.011
12 $\approx 120,000$ 0.232 30 0.014
13 $\approx 205,000$ 0.235 25 0.016
14 $\approx 330,000$ 0.236 20 0.016
QML Amplitude Encoding None 65,825 0.228 55 0.008
Table 11: The total number of trainable parameters and key outputs of the XGBoost with different tree depths. The horizontal line indicates the cases where the XGBoost achieves a lower average RMSE (above the horizontal line) than our QML with Amplitude Encoding model.
Figure 13: The average test RMSE calculated over four cross-validation folds of the XGBoost model with different tree depths. The dots indicate the corresponding epoch of the minimum RMSE achieved (which essentially indicates when the model converges).
Table 11 and Figure 13 offer several key findings. First, increasing tree depth leads to improved performance up to a point, specifically, depth six yields the lowest average RMSE. Beyond this, deeper trees result in no further accuracy gains and even degrade performance, likely due to overfitting. However, while shallower trees mitigate overfitting, they require a larger number of tree estimators to achieve convergence. For example, with depth six, XGBoost required around 270 trees. This increases model complexity and training time, reducing the overall efficiency of the approach. In addition, compared to XGBoost models with depth greater than nine, our QML model achieves better generalization performance (lower RMSE) while maintaining the same computational cost. Notably, the QML model also achieves the lowest standard deviation across test runs. This suggests that, despite XGBoost’s lower RMSE with depth lower than nine, the difference may not be statistically significant due to the higher variability in XGBoost’s results (see Section 5.4).
Appendix B Early stage results on QPU and Noisy Simulator
B.1 Preliminary results on real QPU Rigetti-Ankaa3 of the QML with Angle encoding
In this appendix, we present preliminary results obtained using the proposed QML model on both a noisy quantum simulator and real quantum hardware. Before discussing the results, it is important to note that due to significant resource constraints, we trained our model on a relatively small dataset comprising 100 samples and 32 features, which we trained for just 13 epochs on the quantum hardware. This dataset was split into a single fold, with 75% used for training and 25% for testing.
Due to hardware compatibility limitations, only the QML model employing Angle Encoding was tested on a real quantum processor (see Figure 5(a)).
The real QPU used in this study was Rigetti’s Ankaa-3, a universal, gate-based quantum processor built with tunable superconducting qubits. Detailed specifications of the QPU can be found in [81] and are summarized in Table 12.
For comparison purposes, we also implemented a noisy simulation of the quantum model. The noise model was designed based on the specifications of the Rigetti Ankaa-3 QPU [81], as outlined in Table 12. The parameters used in the noisy simulation are reported in Table 13.
Hardware Specifications
Technology $n$ Qubits Native Gates Gate Time Gradient
[…]
Noise Specifications
Median T1 Median T2 Median f1QRB Median fISWAP Median fRO
[…]
Table 12: Specifications of the Rigetti Ankaa3 Quantum hardware [81]. The T1 inidactes the relaxation time, so it indicates the time one qubit is able to keep is excited status corresponding to |$\ket{1}$ before collapsing in the ground state |$\ket{0}$. T2 inidactes the dephasing time, the time one qubit is able to keep is pure superposition states before, because of the interaction with the environment, to become a random mixture of quantum states. The Median f1QRB indicates the median fidelity of single-qubit gates measured using Randomized Benchmarking (RB), while the Median fISWAP for the two-qubit native gates. The Median fRO indicate the median fidelity due to read out operation of one qubit.
Prob. 1Q Depolarizing Channel Prob. 2Q Depolarizing Channel Prob. Amplitude Damping Prob. Dephasing Prob. Read-Out Error
0.09% 1.42% 0.023% 0.037% 5.1%
Table 13: Noisy parameter chosen to simulate the real QPU Rigetti Ankaa-3. These parameters are calculated accordingly to the value in Table 12.
Although our results are still preliminary and do not support a definitive conclusion, Figure 14(a) demonstrates the promising noise resilience of the QML model with Angle Encoding: during the first 13 epochs, both the noisy simulation and the execution on a real QPU outperform the classical FNN in terms of RMSE value.
(a) Angle Encoding on Real QPU
(b) Amplitude Encoding on Noisy Simulator.
Figure 14: (Left panel) The average test RMSE over ten different experiments of the FNN, the noisy QML model with Angle Encoding, and one single experiment on a real QPU. The models have been trained on a reduced dataset with one hundred observations and 32 features. (Right panel) The test RMSE of the noisy QML model with Amplitude Encoding, the noiseless corresponding model trained on the entire dataset.
B.2 Preliminary results on noisy simulator of QML with Amplitude encoding
Using the same noisy model described in Table 13, we tested the full QML model with Amplitude Encoding, as shown in Figure 4. Due to the high computational cost of simulating noise on classical hardware, some limitations apply. We were able to train the QML with Amplitude Encoding on the full dataset for one hundred epochs, but only a single train/test split (75%/25%) was used at this stage. Figure 14(b) indicates that the QML model with Amplitude Encoding also exhibits good noise resilience, showing only minor deviations from the noiseless model.